OpenStudy (anonymous):

find the values of s for which the quadratic equation x^2-2(s-3)x+(s^2-4s-5)=0 has real roots

hartnn (hartnn):

for real roots, the discriminant \(b^2-4ac \ge 0\) can you find a,b,c for this ?

OpenStudy (anonymous):

how would you find the values of a, b and c?

hartnn (hartnn):

compare \(x^2-2(s-3)x+(s^2-4s-5)=0 \) with \(ax^2+bx+c=0\)

OpenStudy (anonymous):

i dont understand, would the b value be 2(s-3)x?

hartnn (hartnn):

'b' is the co-efficient of x so, b = -2(s-3)

OpenStudy (anonymous):

oh, so what about for a then? would be 1 because its x^2?

OpenStudy (anonymous):

or would it be 0?

hartnn (hartnn):

ax^2 ---->1x^2 --->a=1

OpenStudy (anonymous):

oh yeah